Consider all sample siza 5 fro this population 2 5 6 8 10 12 13 compute the mean u and standard deviation ó of the population


1
Expert's answer
2022-03-19T08:54:20-0400

"EX=\frac{2+5+6+8+10+12+13}{7}=8\\EX^2=\frac{2^2+5^2+6^2+8^2+10^2+12^2+13^2}{7}=77.4286\\DX=EX^2-\left( EX \right) ^2=77.4286-8^2=13.4286\\sX=\sqrt{DX}=\sqrt{13.4286}=3.66451"